PROBLEM 16
~~~~~~~~~~~~~~~~
It takes Daniel ten hours to oil the lanes in a bowling alley. Ndiba can oil the same lanes in 9 hours. If they worked together how long would it take them?

Answer by rwm(914)

x/10+x/9=1
We know it should be a little less than 5 hrs together. Why?
9x+10x=90
19x=90
x=90/19 =4.74 hrs.